# Nonparametric Jackknife Instrumental Variable Estimation and Confounding Robust Surrogate Indices

### v1

@@ B.4.1 First empirical process term
$\langle\Pi(h-h_{0}),(\widehat{\Pi}_{1}-\Pi)(h-h_{0})\rangle$ (53)
$=\frac{2}{N}\sum_{i:V_{i}=1}[\Pi(h-h_{0})](A_{i})(h-h_{0})(S_{i})-E[[\Pi(h-h_{0})](A_{i})(h-h_{0})(S_{i})],$ (54)
@@
$\langle\Pi(h-h_{0}),(\widehat{\Pi}_{1}-\Pi)(h-h_{0})\rangle$ (55)
$\leq\left\lVert[\Pi(h-h_{0})]\times(h-h_{0})\right\rVert\bar{\delta}_{N}+\bar{\delta}_{N}^{2}$ (56)
$\leq\left\lVert[\Pi(h-h_{0})]\right\rVert\bar{\delta}_{N}+\bar{\delta}_{N}^{2}.$ (57)
@@ B.4.2 Second empirical process term
Let $\breve{\delta}_{N}$ be an upper bound on the critical radius of $\Pi\mathcal{G}$.
@@
$\langle\Pi g,\widehat{\Pi}_{0}\eta\rangle=$ $\frac{1}{N/2}\sum_{i:V_{i}=0}[\Pi g](A_{i})\eta_{i}-\mathbb{E}[[\Pi g](A_{i})\eta_{i}],$ (58)
@@
$\langle\Pi g,\widehat{\Pi}_{0}\eta\rangle=O\left(\sqrt{\log N}\left(\breve{\delta}_{N}\|\Pi g\|+\breve{\delta}_{N}^{2}\right)\right).$ (60)
@@ B.5 Empirical and population risk difference for the primary nuisance
$\langle(\widehat{\Pi}_{0}-\Pi)g,(\widehat{\Pi}_{1}-\Pi)g\rangle$ $\widetilde{\delta}_{n}$, $\widehat{\delta}_{N}$ $\mathcal{G}^{0}$, $(\widehat{\Pi}_{0}\mathcal{G}^{0})\times\mathcal{G}^{0}$ $\widehat{\delta}_{N}\widetilde{\delta}_{n}\|g^{0}\|^{2}+\widehat{\delta}_{N}\widetilde{\delta}_{n}^{2}\|g^{0}\|+\widehat{\delta}_{N}^{2}$
$\langle\Pi g,(\widehat{\Pi}_{v}-\Pi)g\rangle$ $\bar{\delta}_{N}$ $(\Pi\mathcal{G})\times\mathcal{G}$ $\bar{\delta}_{N}\|\Pi g\|+\bar{\delta}_{N}^{2}$
$\langle\Pi g,\widehat{\Pi}_{v}\eta\rangle$ $\breve{\delta}_{N}$ $\Pi\mathcal{G}$ $\breve{\delta}_{N}\|\Pi g\|+\breve{\delta}_{N}^{2}$
$\langle(\widehat{\Pi}_{1-v}-\Pi)g,(\widehat{\Pi}_{v}-\Pi)\eta\rangle$ $\widetilde{\delta}_{N}$ $\mathcal{G}^{0}$ $n^{-1/2}(\widetilde{\delta}_{N}\|g^{0}\|+\widetilde{\delta}_{N}^{2})$
@@
$\widehat{R}_{1,0}(h)-R_{1}(h)=O\left(\delta_{N}\|\Pi(h-h_{0})\|+\delta_{N}\widetilde{\delta}_{n}\|h-h_{0}\|^{2}+\delta_{N}n^{-1/2}\|h-h_{0}\|+\delta_{N}^{2}\right).$ (61)
@@ B.6 Weak norm convergence rate for the primary nuisance
Suppose the assumptions of 11 hold, and that $\delta_{N}\widetilde{\delta}_{n}=O(\lambda)$. Then it holds with probability $1-O(\zeta)$ that
$\|\Pi(\widehat{h}-h_{0})\|^{2}=O(\lambda+\delta_{N}\delta_{n}).$ (62)
@@
$\|\Pi(\widehat{h}-h_{0})\|^{2}$ (63)
$\leq$ $\|\Pi(h^{*}-h_{0})\|^{2}$ (64)
$+O\left(\delta_{N}(\|\Pi(\widehat{h}-h^{*})\|+\|\Pi(h^{*}-h_{0})\|)\right.$ (65)
@@
$\leq$ $2\|\Pi(h_{0}-h^{*})\|^{2}$ (70)
$+O(\delta_{N}\|\Pi(\widehat{h}-h^{*})\|+\delta_{N}\delta_{n}\|\widehat{h}-h^{*}\|^{2}+\delta_{N}n^{-1/2}\|\widehat{h}-h^{*}\|+\delta_{N}\delta_{n})$ (71)
@@
From an analysis of the second-order polynomial $t\mapsto\|\Pi(h^{*}+t(\widehat{h}-h^{*})-h_{0})\|^{2}+\|h^{*}+t(\widehat{h}-h^{*})-h_{0}\|^{2}$,
$\|\Pi(\widehat{h}-h^{*})\|^{2}+\lambda\|\widehat{h}-h^{*}\|^{2}\leq\|\Pi(\widehat{h}-h_{0})\|^{2}-\|\Pi(h^{*}-h_{0})\|^{2}-\lambda(\|h^{*}\|^{2}-\|\widehat{h}\|^{2}).$ (73)
@@
$\|\Pi(\widehat{h}-h^{*})\|^{2}+\lambda\|\widehat{h}-h^{*}\|^{2}$ (74)
$\leq$ $\|\Pi(h^{*}-h_{0})\|^{2}$ (75)
$+O(\delta_{N}\|\Pi(h^{*}-h_{0})\|+\delta_{N}\delta_{n}\|\widehat{h}-h^{*}\|^{2}+\delta_{N}\delta_{n}^{2}\|\widehat{h}-h^{*}\|+\delta_{N}\delta_{n}+\lambda\delta_{N}\|\widehat{h}-h^{*}\|).$ (76)
Applying the AM-GM inequality to $\delta_{N}\|\Pi(h^{*}-h_{0})\|$ and using that $\delta_{N}\delta_{n}=O(\lambda)$ and that $\|\Pi(h^{*}-h_{0})\|^{2}=O(\lambda)$ yields that
$\frac{1}{2}\|\Pi(\widehat{h}-h^{*})\|^{2}+\frac{\lambda}{2}\|\widehat{h}-h^{*}\|^{2}=O(\lambda+\delta_{N}\delta_{n}),$ (77)
@@ B.7 Consistency in strong norm of the primary nuisance estimator
$\frac{1}{2}\|\Pi(\widehat{h}-h^{\dagger})\|^{2}\leq$ $R_{1}(\widehat{h})-R_{1}(h^{\dagger})$ (78)
$\leq$ $\widehat{R}_{1,0}(\widehat{h})-\widehat{R}_{1}(h^{\dagger})+O(\delta_{N}\|\Pi(\widehat{h}-h^{\dagger})\|+\delta_{N}\widetilde{\delta}_{n})$ (79)
$\leq$ $O(\delta_{N}\|\Pi(\widehat{h}-h^{\dagger})\|+\delta_{N}\widetilde{\delta}_{n})-\lambda(\|\widehat{h}\|_{2,N}^{2}-\|h^{\dagger}\|^{2}_{2,N}).$ (80)
@@ C.1 Regularization bias under source
$\|\xi-\xi_{0}\|=O(\|w_{0}\|^{2}\lambda^{\beta})\qquad\text{and}\qquad\|\Pi(\xi-\xi_{0})\|^{2}=O(\|w_{0}\|^{2}\lambda^{1+\beta}).$ (82)
@@ C.2 Decomposition of the equicontinuity term
$=$ $\frac{1}{2}\langle(\widehat{\Pi}_{0}-\Pi)(\xi-\xi_{0}),(\widehat{\Pi}_{1}-\Pi)(\xi-\xi_{0})\rangle$ (86)
$+\sum_{v=0,1}\left\{\langle\Pi(\xi-\xi_{0}),(\widehat{\Pi}_{v}-\Pi)(\xi-\xi_{0})\rangle\right.$ (87)
$\qquad+\langle(\widehat{\Pi}_{v}-\Pi)\xi_{0},(\widehat{\Pi}_{1-v}-\Pi)(\xi-\xi_{0})\rangle$ (88)
$\left.\qquad+\langle\Pi\xi_{0},(\widehat{\Pi}_{v}-\Pi)(\xi-\xi_{0})\rangle\right\}.$ (89)

### v2

@@ B.4.1 First empirical process term
$\langle\mathcal{T}_{K}(h-h_{0}),(\widehat{\mathcal{T}}_{K,1}-\mathcal{T}_{K})(h-h_{0})\rangle$ (72)
$=\frac{2}{N}\sum_{i:V_{i}=1}[\mathcal{T}_{K}(h-h_{0})](A_{i})(h-h_{0})(S_{i})-E[[\mathcal{T}_{K}(h-h_{0})](A_{i})(h-h_{0})(S_{i})],$ (73)
@@
$\langle\mathcal{T}_{K}(h-h_{0}),(\widehat{\mathcal{T}}_{K,1}-\mathcal{T}_{K})(h-h_{0})\rangle$ (74)
$\leq\left\lVert[\mathcal{T}_{K}(h-h_{0})]\times(h-h_{0})\right\rVert\bar{\delta}_{N}+\bar{\delta}_{N}^{2}$ (75)
$\leq\left\lVert[\mathcal{T}_{K}(h-h_{0})]\right\rVert\bar{\delta}_{N}+\bar{\delta}_{N}^{2}.$ (76)
@@ B.4.2 Second empirical process term
Let $\breve{\delta}_{N}$ be an upper bound on the critical radius of $\mathcal{T}_{K}\mathcal{G}$.
@@
$\langle\mathcal{T}_{K}g,\widehat{\mathcal{T}}_{K,0}\eta\rangle=$ $\frac{1}{N/2}\sum_{i:V_{i}=0}[\mathcal{T}_{K}g](A_{i})\eta_{i}-\mathbb{E}[[\mathcal{T}_{K}g](A_{i})\eta_{i}],$ (77)
@@
$\langle\mathcal{T}_{K}g,\widehat{\mathcal{T}}_{K,0}\eta\rangle=O\left(\sqrt{\log N}\left(\breve{\delta}_{N}\|\mathcal{T}_{K}g\|+\breve{\delta}_{N}^{2}\right)\right).$ (79)
@@ B.5 Empirical and population risk difference for the primary nuisance
$\langle(\widehat{\mathcal{T}}_{K,0}-\mathcal{T}_{K})g,(\widehat{\mathcal{T}}_{K,1}-\mathcal{T}_{K})g\rangle$ $\widetilde{\delta}_{n}$, $\widehat{\delta}_{N}$ $\mathcal{G}^{0}$, $(\widehat{\mathcal{T}}_{K,0}\mathcal{G}^{0})\times\mathcal{G}^{0}$ $\widehat{\delta}_{N}\widetilde{\delta}_{n}\|g^{0}\|^{2}+\widehat{\delta}_{N}\widetilde{\delta}_{n}^{2}\|g^{0}\|+\widehat{\delta}_{N}^{2}$
$\langle\mathcal{T}_{K}g,(\widehat{\mathcal{T}}_{K,v}-\mathcal{T}_{K})g\rangle$ $\bar{\delta}_{N}$ $(\mathcal{T}_{K}\mathcal{G})\times\mathcal{G}$ $\bar{\delta}_{N}\|\mathcal{T}_{K}g\|+\bar{\delta}_{N}^{2}$
$\langle\mathcal{T}_{K}g,\widehat{\mathcal{T}}_{K,v}\eta\rangle$ $\breve{\delta}_{N}$ $\mathcal{T}_{K}\mathcal{G}$ $\breve{\delta}_{N}\|\mathcal{T}_{K}g\|+\breve{\delta}_{N}^{2}$
$\langle(\widehat{\mathcal{T}}_{K,1-v}-\mathcal{T}_{K})g,(\widehat{\mathcal{T}}_{K,v}-\mathcal{T}_{K})\eta\rangle$ $\widetilde{\delta}_{N}$ $\mathcal{G}^{0}$ $n^{-1/2}(\widetilde{\delta}_{N}\|g^{0}\|+\widetilde{\delta}_{N}^{2})$
@@
$\widehat{R}_{1,0}(h)-R_{1}(h)=O\left(\delta_{N}\|\mathcal{T}_{K}(h-h_{0})\|+\delta_{N}\widetilde{\delta}_{n}\|h-h_{0}\|^{2}+\delta_{N}n^{-1/2}\|h-h_{0}\|+\delta_{N}^{2}\right).$ (80)
@@ B.6 Weak norm convergence rate for the primary nuisance
Suppose the assumptions of 12 hold, and that $\delta_{N}\widetilde{\delta}_{n}=O(\lambda)$. Then it holds with probability $1-O(\zeta)$ that
$\|\mathcal{T}_{K}(\widehat{h}-h_{0})\|^{2}=O(\lambda+\delta_{N}\delta_{n}).$ (81)
@@
$\|\mathcal{T}_{K}(\widehat{h}-h_{0})\|^{2}$ (82)
$\leq$ $\|\mathcal{T}_{K}(h^{*}-h_{0})\|^{2}$ (83)
$+O\left(\delta_{N}(\|\mathcal{T}_{K}(\widehat{h}-h^{*})\|+\|\mathcal{T}_{K}(h^{*}-h_{0})\|)\right.$ (84)
@@
$\leq$ $2\|\mathcal{T}_{K}(h_{0}-h^{*})\|^{2}$ (89)
$+O(\delta_{N}\|\mathcal{T}_{K}(\widehat{h}-h^{*})\|+\delta_{N}\delta_{n}\|\widehat{h}-h^{*}\|^{2}+\delta_{N}n^{-1/2}\|\widehat{h}-h^{*}\|+\delta_{N}\delta_{n})$ (90)
@@
From an analysis of the second-order polynomial $t\mapsto\|\mathcal{T}_{K}(h^{*}+t(\widehat{h}-h^{*})-h_{0})\|^{2}+\|h^{*}+t(\widehat{h}-h^{*})-h_{0}\|^{2}$,
$\|\mathcal{T}_{K}(\widehat{h}-h^{*})\|^{2}+\lambda\|\widehat{h}-h^{*}\|^{2}\leq\|\mathcal{T}_{K}(\widehat{h}-h_{0})\|^{2}-\|\mathcal{T}_{K}(h^{*}-h_{0})\|^{2}-\lambda(\|h^{*}\|^{2}-\|\widehat{h}\|^{2}).$ (92)
@@
$\|\mathcal{T}_{K}(\widehat{h}-h^{*})\|^{2}+\lambda\|\widehat{h}-h^{*}\|^{2}$ (93)
$\leq$ $\|\mathcal{T}_{K}(h^{*}-h_{0})\|^{2}$ (94)
$+O(\delta_{N}\|\mathcal{T}_{K}(h^{*}-h_{0})\|+\delta_{N}\delta_{n}\|\widehat{h}-h^{*}\|^{2}+\delta_{N}\delta_{n}^{2}\|\widehat{h}-h^{*}\|+\delta_{N}\delta_{n}+\lambda\delta_{N}\|\widehat{h}-h^{*}\|).$ (95)
Applying the AM-GM inequality to $\delta_{N}\|\mathcal{T}_{K}(h^{*}-h_{0})\|$ and using that $\delta_{N}\delta_{n}=O(\lambda)$ and that $\|\mathcal{T}_{K}(h^{*}-h_{0})\|^{2}=O(\lambda)$ yields that
$\frac{1}{2}\|\mathcal{T}_{K}(\widehat{h}-h^{*})\|^{2}+\frac{\lambda}{2}\|\widehat{h}-h^{*}\|^{2}=O(\lambda+\delta_{N}\delta_{n}),$ (96)
@@ B.7 Consistency in strong norm of the primary nuisance estimator
$\frac{1}{2}\|\mathcal{T}_{K}(\widehat{h}-h^{\dagger})\|^{2}\leq$ $R_{1}(\widehat{h})-R_{1}(h^{\dagger})$ (97)
$\leq$ $\widehat{R}_{1,0}(\widehat{h})-\widehat{R}_{1}(h^{\dagger})+O(\delta_{N}\|\mathcal{T}_{K}(\widehat{h}-h^{\dagger})\|+\delta_{N}\widetilde{\delta}_{n})$ (98)
$\leq$ $O(\delta_{N}\|\mathcal{T}_{K}(\widehat{h}-h^{\dagger})\|+\delta_{N}\widetilde{\delta}_{n})-\lambda(\|\widehat{h}\|_{2,N}^{2}-\|h^{\dagger}\|^{2}_{2,N}).$ (99)
@@ C.1 Regularization bias under source
$\|\xi-\xi_{0}\|=O(\|w_{0}\|^{2}\lambda^{\beta})\qquad\text{and}\qquad\|\mathcal{T}_{K}(\xi-\xi_{0})\|^{2}=O(\|w_{0}\|^{2}\lambda^{1+\beta}).$ (101)
@@ C.2 Decomposition of the equicontinuity term
$=$ $\frac{1}{2}\langle(\widehat{\mathcal{T}}_{K,0}-\mathcal{T}_{K})(\xi-\xi_{0}),(\widehat{\mathcal{T}}_{K,1}-\mathcal{T}_{K})(\xi-\xi_{0})\rangle$ (105)
$+\sum_{v=0,1}\left\{\langle\mathcal{T}_{K}(\xi-\xi_{0}),(\widehat{\mathcal{T}}_{K,v}-\Pi)(\xi-\xi_{0})\rangle\right.$ (106)
$\qquad+\langle(\widehat{\mathcal{T}}_{K,v}-\Pi)\xi_{0},(\widehat{\mathcal{T}}_{K,1-v}-\Pi)(\xi-\xi_{0})\rangle$ (107)
$\left.\qquad+\langle\Pi\xi_{0},(\widehat{\mathcal{T}}_{K,v}-\Pi)(\xi-\xi_{0})\rangle\right\}.$ (108)